The production function in an economy is Y equals A left parenthesis L plus square root of K right parenthesis where A is productivity, L is the number of workers, and K is the amount of capital stock. Suppose the current level of A is 1, K is 2500 and the labor market clears when 900 workers are employed. Money demand is M over P equals 0.5 Y minus 250 left parenthesis r plus straight pi to the power of straight e right parenthesis The expected rate of inflation is 0.02, the nominal money supply is 9100 and the price level is 20. Desired consumption is C to the power of d equals 262 plus 0.4 Y minus 200 r and desired investment is I to the power of d equals 265 minus 250 r Government purchase spending is G=70. 2) [4 points] What is the real interest rate, r, that clears the asset market when Y=1000? when Y=950?
Answer #1

☆ -0.5Y - 250 (v +10. = 0.02 ; M = 9100 y = 10oo. ; P=20 109100.05 (100) - 250 (1 +0.02) 20 455 50 - 2500-5 455 = 495-2500 25( M = 505) - 250 ( + ). 0 . 02 = ae_ P مو = Y = 950 20 = ) 250 - ( میه) 6 . 5 - صورة ( = (10 . 0 2o 4ss : 435 - 25 - 5 که 250